OpenStudy (anonymous):

Find the quotient and express the answer in scientific notation. 302 (9.1 x 104)

OpenStudy (anonymous):

its 302 divided by 9.1 times 10 to the fourth power :)

OpenStudy (theviper):

\[\Huge{\color{red}{\frac{302}{9.2\times10^4}}}\]

OpenStudy (anonymous):

right! now what do i do?

OpenStudy (shane_b):

Divide it and convert it to scientific notation. What do you get when you divide it?

OpenStudy (theviper):

@Shane_B is right:)

OpenStudy (theviper):

Now @lam22 can do it as I think :)

OpenStudy (anonymous):

I think its 910,000

OpenStudy (shane_b):

Get a new calculator :)

OpenStudy (anonymous):

I wasnt using one....

OpenStudy (theviper):

@mathslover help @lam22 :)

OpenStudy (shane_b):

This is simple division:\[\frac{302^1}{9.2x10^4}\]

OpenStudy (anonymous):

I have no clue what the heck I am doing....

OpenStudy (shane_b):

\[\frac{302}{9.2}=?\]

mathslover (mathslover):

\[\huge{\frac{302}{9.1*10^4}}\] \[\huge{\frac{3020}{91*10^4}}\] \[\huge{\frac{33.1}{10^4}}\] \[\huge{\frac{331}{10^5}}\] \[\huge{331*10^{-5}}\] \[\huge{3.31*10^{-3}}\]

OpenStudy (shane_b):

I was trying to let him do it :P

OpenStudy (shane_b):

There's a mistake in yours btw...it was 9.2. But he should still be able to figure it out.

OpenStudy (theviper):

got it @lam22 ?

mathslover (mathslover):

oh sorry @Shane_B for that... i thought he needed the answer now

OpenStudy (theviper):

yes @mathslover

mathslover (mathslover):

so i typed fastly and did that mistake

mathslover (mathslover):

sorry

OpenStudy (shane_b):

minor mistake...no biggie.

mathslover (mathslover):

well i think @lam22 would be able to do that still

OpenStudy (shane_b):

@lam22: Do you understand the steps?

OpenStudy (theviper):

what a catching power @Shane_B !!

OpenStudy (shane_b):

@mathslover: How do you get the big numbers btw? I'm still learning the syntax

mathslover (mathslover):

easy .. \ .[ huge{\text{hi}}\.} remove dots

OpenStudy (shane_b):

I was trying to get a large fraction bar earlier too but couldn't seem to do it...like when dividing a fraction by a fraction.

OpenStudy (shane_b):

ahh

OpenStudy (theviper):

Type this in equation editor:- Huge{color{red}{I heartsuit Openstudy.}}

OpenStudy (theviper):

\[\Huge{\color{red}{I \heartsuit Openstudy.}}\]

mathslover (mathslover):

\[\Huge{\color{red}{I \heartsuit Openstudy.}.}.\]

OpenStudy (theviper):

woah!!

OpenStudy (shane_b):

\[\Huge{\color{red}{I \heartsuit Openstudy.}}\]Neat...didn't know you could do colors too.

OpenStudy (shane_b):

What do you guys use for a latex reference on all this?

mathslover (mathslover):

@TheViper do u know how to insert annoyingly large fonts ?

OpenStudy (theviper):

Type this:- @Shane_B Huge{color{gold}{star star}color{blue}{I heartsuit Openstudy.}}

OpenStudy (theviper):

\[\Huge{\color{gold}{\star \star}\color{blue}{I \heartsuit Openstudy.}}\]

mathslover (mathslover):

@TheViper do u know how to insert annoyingly large fonts

OpenStudy (shane_b):

Gold stars FTW!

mathslover (mathslover):

U mean Huge and huge have difference

OpenStudy (theviper):

Nahi mujhe bhi thoda samajh nahi aaya:/